OpenStudy (anonymous):

What is the equation of a line passing through (-6, 5) and having a slope of ?

OpenStudy (mosaic):

What is the slope? It is missing in the question.

OpenStudy (anonymous):

1/3

OpenStudy (mosaic):

y = mx + b Slope m = 1/3. Therefore, y = 1/3x + b (-6,5) is a point on the line. Put x = -6 and y = 5 and solve for b.